# FaceSORT: a Multi-Face Tracking Method based on Biometric and Appearance Features

### v1

@@ 6 Conclusion
It this work, a new multi-face tracking method, FaceSORT, is proposed. To mitigate the problem of partially occluded and lateral faces, two different features (i.e., face (biometric) and appearance features) are combined. In the considered scenario, when people move towards a gate/portal (PLUSFiaQ and ChokePoint dataset) FaceSORT clearly outperformed the evaluated state-of-the-art tracker. To get a deeper insight into the proposed method, a comprehensive experimental evaluation and an ablation study are conducted. It is shown that the selection of the face descriptor and the similarity measure, as well as the resulting distribution of genuine and imposter scores are crucial. In general, selecting a low similarity threshold $\theta$ and a high parameter $\lambda$ lead to a high AssPr.

### v2

@@ 6 Conclusion
It this work, a new multi-face tracking method, FaceQSORT, is proposed. To mitigate the problem of partially occluded and lateral faces, two different features (i.e., face (biometric) and appearance features) are combined. In the considered scenario, when people move towards a gate/portal (PLUSFiaQ and ChokePoint dataset) FaceQSORT clearly outperformed the evaluated state-of-the-art tracker. To get a deeper insight into the proposed method, a comprehensive experimental evaluation and an ablation study are conducted. It is shown that the selection of the face descriptor and the similarity measure, as well as the resulting distribution of genuine and imposter scores are crucial. In general, selecting a low similarity threshold $\theta$ and a high parameter $\lambda$ lead to a high AssPr.